# Calabi-Yau Four/Five/Six-folds as $\mathbb{P}_{\textbf{w}}^{n}$ Hypersurfaces: Machine Learning, Approximation, and Generation

### v1

@@ 4.1.2 Symbolic Regression
With the goal of extracting just the first order behaviour for the NN approximate formula, the high parsimony and simple function basis used has limited this regression performance, leading to expected lower performance scores relative to the NNs. However these three independent expression have similar performance, and some similar structure.
The first thing to note is that in each equation there are three summed terms, which are each positive functions of weights. More specifically, each has a term equal to $w_{2}$, and another time either equal or proportional to $w_{1}$. The occurrence of these earlier weights somewhat corroborates the importance of earlier parts of the weight system seen in §4.1.1, however without the $w_{0}$ factor – which may be related to $w_{0}$ having a significantly smaller range. In each case there is one further term involving higher weights, and across these expressions all additional weights do occur in this term. Overall, it is quite surprising how well such simple expressions can perform at predicting the $h^{1,1}$ values, and the simple linear sum behaviour does support there being an approximate linear relationship as observed in Figure 3.3.
@@ 4.2 Classifying CY Property
Focusing on the MCC scores, the architectures struggle most with identifying reflexivity, especially in the CY case. Considering the number of steps required to compute this analytically, via construction of the respective lattice polytope, taking the dual polytope, and then performing many integer checks of the corresponding vertices, this is perhaps not surprising. Conversely, the performance for identifying the IP property is then surprisingly high, which still requires generating the polytope. Therefore it is likely the NNs can approximate the polytopes within their architectures but struggle with the integer checks of vertices – a property notoriously evasive for ML testolin2023neural . Respectively, the NNs can well learn the intradivisibility property, a more direct computation with the weight data, however still with a long list of necessary checks. Finally, and most pleasingly, the CY property can be well learnt for four-fold weight systems, a result observed for three-folds in Berman_2022 . The ability to so successfully predict the existence of further singularity structure in the respective hypersurfaces beyond that of the ambient weighted projective space remains astounding, for a method which is still unclear how to perform directly without using the Landau-Ginzburg string interpretation.
@@ 5.1 Presenting the Formula
$\sum_{p,q}(-1)^{p+q}h^{p,q}u^{p}v^{q}=\frac{1}{uv}\sum_{0\leq l<w_{tot}}\left[\prod_{\tilde{\theta}_{i}(l)=0}\frac{(uv)^{q_{i}}-uv}{1-(uv)^{q_{i}}}\right]_{int}\cdot(-u)^{\operatorname{size}(l)}\left(\frac{v}{u}\right)^{\operatorname{age}(l)}\,.$ (5.4)
It is evident that the polynomial products and divisions within the square brackets are what takes most of the computational resources. Just for reference, we note that for high weights, the software sagemath cannot perform the calculation due to the high number of terms. For this reason, the algorithm had to be hardcoded directly. One way to go about simplifying this formula is to identify for which values of $l$ the terms in the sum contribute the most. Just by empirical observation, we find that the main contributions come from the element of zero age and the ones with maximal size [^16] This can be explicitly verified by looking at examples 4.5 and 4.6 in Batyrev:2020ych .. As we are about to argue in detail, it turns out that including only terms of these types provides a very efficient approximation for the Hodge numbers. It is both very accurate and much faster to implement. Moreover, it bounds the exact results from below. Explicitly, the approximated Hodge numbers $h_{A}^{p,q}$ can be computed as:
$\begin{split}\sum_{p,q}(-1)^{p+q}h_{A}^{p,q}u^{p}v^{q}=\frac{1}{uv}\Bigg{\{}\left[\prod_{i}\frac{(uv)^{q_{i}}-uv}{1-(uv)^{q_{i}}}\right]_{int}+\sum_{\mathrm{size}(l)=n}(-u)^{d}\left(\frac{v}{u}\right)^{\operatorname{age}(l)}\Bigg{\}}\,.\end{split}$ (5.5)
We tested this approximation against the two relevant datasets: the Calabi-Yau four-folds considered in this article and the shorter set of Calabi-Yau three-folds. We start by presenting our findings for the latter case [^17] For completeness, we point out that the approximation also fails to be well-defined for non-Calabi-Yau weight systems in general..
@@ 5.2 Application to Calabi-Yau Three-folds
As just mentioned, (5.4) becomes more and more involved to compute as the weights in the system become larger. Since the dataset for Calabi-Yau manifolds in weighted projective spaces are already ordered according to the sum of the weights, we conveniently divided the $7555$ three-folds’ weight systems (of 5 weights) into $11$ groups, from lowest to highest. The plot in Figure 0(a) shows the mean percentage error of the approximation formula (5.5) for each of those groups, both for $h^{1,1}$ and $h^{1,2}$, showing that the precision increases as we go to higher weights. For reference, we also include a measure for the mean sum of weights in each of the groups $\langle 5/w_{tot}\rangle$. We observe that the average percentage error for large weights is remarkably small, lying somewhere between $3\%$ and $5\%$ for both Hodge numbers.
@@ 5.3 Application to Calabi-Yau Four-folds
We now move to the case of four-folds in weighted projective spaces. This dataset is considerably bigger compared to three-folds, with $1100055$ spaces, and it contains systems with very large weights. A natural consequence is that the computational times are much longer, which makes the advantages of the approximation even more evident. To give a concrete example, we focused on the millionth weight system in the dataset, which reads $[45,74,2460,12792,17876,33173]$. Our implementation of (5.4) takes roughly $40$ hours, while the approximated version (5.5) is computed in $49$ seconds. The ratio between the two is $0.00034$.
Regarding the precision of the approximation, we illustrate it in Figure 2(a), with a similar plot to the one used for three-folds. We omit $h^{1,2}$ from the picture because it can take the value of zero, making the percentage error not well-defined. We sampled randomly and uniformly $20\%$ of the dataset, and then divided it into groups of $2000$ samples. As before, we ordered the group according to the sum of weights, as shown by the gray points. We observe once again that the percentage error gets smaller as the weights become larger, reaching roughly $1\%$ for the samples with largest weights, for all three Hodge numbers.
@@
The plot in Figure 2(b), on the other hand, shows the comparison between the computational resources employed by the exact expression from the Landau-Ginzburg model and our approximation. It is evident from the example just discussed that systems with large weights necessitate a very long compute time for the full formula (5.4). Thus, we collected data within the first half of the dataset, and then extrapolated our findings to the second half, containing very large weights. The red curve provides a good interpolation of the data, and it turns out to give very accurate predictions as well. This can be confirmed by plugging $w_{tot}=66420$, which corresponds to the weight system discussed above, into the expression for the best fit function. The result is $t_{app}/t=0.00035$, which is remarkably close to the actual ratio obtained from the explicit computation.
@@ 6 Higher Weight systems
The approximation has been tested on both three-folds and four-folds. These are the only two existing datasets of Calabi-Yau manifolds built as hypersurfaces in weighted projective spaces. Here, we present the first efforts towards the understanding of these spaces in higher dimensions, by generating a partial dataset of candidate transverse weight systems of 7 weights, and the respective five-folds’ Hodge numbers. We discuss how the approximation can be used to quickly extract information about such a dataset, and we additionally generate a first partial dataset of candidate transverse weight systems of 8 weights, then using (5.5) to construct an approximated list of six-folds’ Hodge numbers.
@@ 6.1 Calabi-Yau Five-folds
Calabi-Yau five-folds appear in a number of dimensional reductions in the literature. For instance, it was found that M-theory compactified on a Calabi-Yau five-fold results in an exotic $\mathcal{N}=2$ supersymmetric quantum mechanics Haupt:2008nu . Moreover, Calabi-Yau five-folds play a role in F-theory, where upon compactification, they provide a way to systematically construct $N=(0,2)$ CFTs Schafer-Nameki:2016cfr ; Tian:2020yex , which may lead to their classification. Therefore, an extended dataset of such Calabi-Yau manifolds would make it possible to explore the landscape of such CFTs. Additionally, in Curio:1998bv , a 3-dimensional string vacua with $\mathcal{N}=1$ supersymmetry has been found, which can be interpreted as a compactification of S-theory on a Calabi-Yau five-fold. Despite their role in the construction of low-dimensional theories, examples of five-folds have not been systematically constructed, until the recent effort in Alawadhi:2023gxa , which focuses on the CICY construction. Here, we present a second in that direction, i.e. we generate, for the first time, a sub-set of Calabi-Yau five-folds obtained as hypersurfaces in $\mathbb{P}^{6}_{\mathrm{w}}$. Specifically, we generate all 7-weight weight systems whose sum of weights $w_{tot}\leq 200$. To efficiently identify those which have the required property to describe a Calabi-Yau (for more details, see §2), we employ a two-step approach. We first systematically search all partitions of each sum of weights up to 200 as weight systems, extracting those which are coprime, IP [^19] Work in Skarke:1996hq showed that transverse weight systems are by necessity IP for any size weight system., and intradivisible; performed with our code functionality. Then, using the approximation as a tool for establishing the Calabi-Yau property. Selecting all the weight systems that are well-defined with respect to (5.5), or equivalently all those such that the polynomial division $\left[\prod_{i}\frac{(uv)^{q_{i}}-uv}{1-(uv)^{q_{i}}}\right]$ gives no reminder. The candidate weight systems identified this way were then all checked with respect to the exact formula (5.4), and all turned out to be well-defined, yielding the full exact Hodge diamond. To provide confidence in the generated data, two non-trivial checks on the cohomological data were performed. First, computing the Euler number from the weights alone with (2.11), and then verifying it agrees with the identity
@@
As it can be guessed by looking at $\left\langle h^{1,1}\right\rangle$ and $\left\langle h^{1,4}\right\rangle$, the subset of spaces considered here does not show mirror symmetry. This is due to the fact that we only restricted ourselves to a small sum of weights, whose mirror-symmetric pairs lie in the large-weights regime. We expect the cohomological data in that regime to be practically inaccessible due to the large computational times associated with (5.4). For this reason, we believe that the presented approximation of §5, which proved to be extremely accurate for large weights in the four-folds investigation, could be a key tool for attempting such a task. Moreover, we also expect the list of all possible Calabi-Yau 7-weight weight systems to be astronomical in size. Once again, the truncated formula (5.5) provides a quickly computable tight lower bound for the Hodge numbers of all those yet undiscovered manifolds.
@@ 6.2 Calabi-Yau Six-folds
While their role in physics is marginal (they could only be employed for compactifications of S-theory), Calabi-Yau six-folds have their own relevance directly within mathematics. These spaces could provide additional information about the - still very mysterious to this day - landscape of Calabi-Yau geometries, and they are the second non-trivial family of Calabi-Yau manifolds in even complex dimensions. Their construction as hypersurfaces of weighted projective spaces involves using 8-weight weight systems, which are both more numerous and effectively infeasible to run bulk computation of exact topological parameters. For these reasons, we find the truncated approximation formula to be especially pertinent, allowing computation of approximated Hodge values for all the generated candidate transverse weight systems with $w_{tot}\leq 200$. Once again, we first identify the IP intradivisible systems, and then select the ones which are well-defined with respect to (5.5), numbering 1482022 candidate transverse weight systems of 8 weights (accessible at this work’s respective GitHub in the same format as for the five-folds). From there the Euler number was also computed exactly with the less computationally-intensive direct formula from weights (2.11). A few examples are reported in Table 6.2.
@@
Similarly to before, the first manifold in Table 6.2 is nothing but the Calabi-Yau six-fold defined by a degree eight polynomial in $\mathbb{P}^{7}$. We find that our results, despite coming from the truncated formula, exactly match the exact Hodge numbers, which can be found in Dumachev2015CompleteIC .
@@ 7 Summary & Outlook
While all of the above was performed using concrete analytic algorithms, statistical machine learning techniques were also applied both to the dataset of Calabi-Yau four-folds and to the partitioned set of more general weight systems. Regarding the former, a fully connected regressor network was shown to predict the cohomological Hodge data, and the Euler number, from the system weights. We found particularly good results, with $R^{2}\sim 0.91$, for $h^{1,1}$, on the whole dataset. For the other invariants, we observed very different results for systems with small weights as opposed to systems with large weights. For instance, $h^{1,3}$ and $h^{2,2}$ showed results with $R^{2}>0.90$ for the half of the dataset containing lower weights, while the accuracy dropped significantly for the other half. These three numbers provide sufficient information to determine the full Hodge diamond, however results were also reported associated with $h^{1,2}$, which had a poor performances since it is zero $48\%$ of the times, and $\chi$, which showed similar results to $h^{1,3}$ and $h^{2,2}$.
@@
Motivated by the strong performances of the neural networks, and inspired by the interpretability of the gradient saliency analysis and symbolic regression, we explored a simpler truncated version of the formula coming from the Landau-Ginzburg model used for calculating the Calabi-Yau Hodge numbers from the ambient $\mathbb{P}^{n}_{\textbf{w}}$’s weight system. This approximation drastically reduces the number of terms involved in the computation, making it easier to study analytically, and substantially faster to compute numerically. Its main features are: it provides a tight lower bound for the Hodge numbers; it is especially accurate for systems with large weights (average accuracy of $99\%$ for the $10000$ systems with largest weights); it is dramatically faster than the exact formula (up to $10^{4}$ times quicker); it reproduces the observed linear clustering behaviour for large weights.
Finally, motivated by the speed improvements available from this approximation, transverse weight systems (satisfying the necessary intradivisible and IP properties, and well-defined with respect to both the approximation and exact Landau-Ginzburg formula) were generated for a sum of weights $w_{tot}\leq 200$, for systems of 7-weights producing Calabi-Yau five-folds. Additionally, where the exact Landau-Ginzburg formula compute time was infeasible for systems of 8-weights, a complementary dataset of candidate transverse weight systems (satisfying the necessary intradivisible and IP properties, and well-defined with respect to just the approximation) was generated, again for a sum of weights $w_{tot}\leq 200$, leading to candidate Calabi-Yau six-folds. Some preliminary analysis of this data, and the respectively computed topological properties is provided, with a thorough analysis, and its full generation for $w_{tot}>200$, left to future work.
These datasets, the respective code for analysis and ML, and an example notebook illustrating functionality to check intradivisibility, compute Euler number, and compute exact and approximate Hodge numbers of an input weight system of any size is all available at this work’s respective GitHub repository.

### v2

@@ 4.1.2 Symbolic Regression
With the goal of extracting just the first order behaviour for the NN approximate formula, the high parsimony and simple function basis used has limited this regression performance, leading to expected lower performance scores relative to the NNs. Generalisation to a broader basis with lower parsimony can provide expressions with higher performance, well demonstrated by a run using the full gplearn basis
{+,,×,÷$\big{\{}+,-,\times,\div,-(\cdot),\sqrt{\cdot},\frac{1}{\cdot},|\cdot|,\text{log}(\cdot),\text{max}(\cdot),\text{min}(\cdot),\text{sin}(\cdot),\text{cos}(\cdot),\text{tan}(\cdot)\big{\}}\;,$ (4.3)
producing the expression
$\sqrt{w_{3}+3\sqrt{w_{4}}+\text{min}\bigg{(}\frac{w_{4}}{\text{log}(\text{cos}(w_{0})-|w_{1}|)},2w_{1}+\text{min}\bigg{(}(w_{3}-w_{0}),...}$ (4.4)
$\overline{\sqrt{w_{2}\text{min}\bigg{(}\sqrt{w_{2}^{2}\text{log}(\text{cos}(w_{0})-|w_{1}|)},\text{min}\bigg{(}(w_{3}-w_{0}),\text{min}\bigg{(}\sqrt{w_{2}w_{3}},\sqrt{\frac{w_{2}^{2}w_{5}}{w_{3}-w_{0}}}\bigg{)}\bigg{)}\bigg{)}}\bigg{)}\bigg{)}}\;,$
with $R^{2}$ score $0.896$. However due to the far higher equation complexity and this relatively minimal increase in performance, the lack of interpretability puts motivation on consideration of the initially specified simple basis, with candidate expressions quoted in Table 4.5. Where these three independent expressions have similar performance, and some similar structure.
The first thing to note is that in each equation there are three summed terms, which are each positive functions of weights. More specifically, each has a term equal to $w_{2}$, and another term either equal or proportional to $w_{1}$. The occurrence of these earlier weights somewhat corroborates the importance of earlier parts of the weight system seen in §4.1.1, however without the $w_{0}$ factor – which may be related to $w_{0}$ having a significantly smaller range. In each case there is one further term involving higher weights, and across these expressions all additional weights do occur in this term. Overall, it is quite surprising how well such simple expressions can perform at predicting the $h^{1,1}$ values, and the simple linear sum behaviour does support there being an approximate linear relationship as observed in Figure 3.3.
@@ 4.2 Classifying CY Property
Focusing on the MCC scores, the architectures struggle most with identifying reflexivity, especially in the CY case. Considering the number of steps required to compute this analytically, via construction of the respective lattice polytope, taking the dual polytope, and then performing many integer checks of the corresponding vertices, this is perhaps not surprising. Conversely, the performance for identifying the IP property is then surprisingly high, which still requires generating the polytope. Therefore it is likely the NNs can approximate the polytopes within their architectures but struggle with the integer checks of vertices – a property notoriously evasive for ML testolin2023neural . Respectively, the NNs can well learn the intradivisibility property, a more direct computation with the weight data, however still with a number of necessary checks. Finally, and most pleasingly, the CY property can be well learnt for four-fold weight systems, a result observed for three-folds in Berman_2022 . The ability to so successfully predict the existence of further singularity structure in the respective hypersurfaces beyond that of the ambient weighted projective space remains astounding, for a method which is still unclear how to perform directly without using the Landau-Ginzburg string interpretation.
@@ 5.1 Presenting the Formula
$\sum_{p,q}(-1)^{p+q}h^{p,q}u^{p}v^{q}=\frac{1}{uv}\sum_{0\leq l<w_{tot}}\left[\prod_{\tilde{\theta}_{i}(l)=0}\frac{(uv)^{q_{i}}-uv}{1-(uv)^{q_{i}}}\right]_{int}(-u)^{\operatorname{size}(l)}\left(\frac{v}{u}\right)^{\operatorname{age}(l)}\,.$ (5.4)
It is evident that the polynomial products and divisions within the square brackets are what takes most of the computational resources. Just for reference, we note that for high weights, the software sagemath cannot perform the calculation due to the high number of terms. For this reason, the algorithm had to be hardcoded directly. One way to go about simplifying this formula is to identify for which values of $l$ the terms in the sum contribute the most. Just by empirical observation, we find that the main contributions come from the element of zero age and the ones with maximal size [^16] This is well exemplified by looking at examples 4.5 and 4.6 in Batyrev:2020ych .. As we are about to argue in detail, it turns out that including only terms of these types provides a very efficient approximation for the Hodge numbers. It is both very accurate and much faster to implement. Moreover, it bounds the exact results from below. Explicitly, the approximated Hodge numbers $h_{A}^{p,q}$ can be computed as:
$\begin{split}\sum_{p,q}(-1)^{p+q}h_{A}^{p,q}u^{p}v^{q}=\frac{1}{uv}\Bigg{\{}\left[\prod_{i}\frac{(uv)^{q_{i}}-uv}{1-(uv)^{q_{i}}}\right]_{int}+\sum_{\mathrm{size}(l)=n}(-u)^{w_{tot}}\left(\frac{v}{u}\right)^{\operatorname{age}(l)}\Bigg{\}}\,.\end{split}$ (5.5)
We tested this approximation against the two relevant datasets: the Calabi-Yau four-folds considered in this article and the smaller set of Calabi-Yau three-folds. We start by presenting our findings for the latter case [^17] For completeness, we point out that the approximation also fails to be well-defined for non-Calabi-Yau weight systems in general..
@@ 5.2 Application to Calabi-Yau Three-folds
As just mentioned, (5.4) becomes more and more involved to compute as the weights in the system become larger. Since the dataset for Calabi-Yau manifolds in weighted projective spaces are already ordered according to the sum of the weights, we conveniently divided the $7555$ three-folds’ weight systems (of 5 weights) into $11$ groups, from lowest to highest. The plot in Figure 1(a) shows the mean percentage error of the approximation formula (5.5) for each of those groups, both for $h^{1,1}$ and $h^{1,2}$, showing that the approximation becomes more precise as we go to higher weights. For reference, we also include a measure for the mean sum of weights in each of the groups $\langle 5/w_{tot}\rangle$. We observe that the average percentage error for large weights is remarkably small, lying somewhere between $3\%$ and $5\%$ for both Hodge numbers.
@@ 5.3 Application to Calabi-Yau Four-folds
We now move to the case of four-folds in weighted projective spaces. This dataset is considerably bigger compared to three-folds, with $1100055$ spaces, and it contains systems with very large weights. A natural consequence is that the computational times are much longer, which makes the advantages of the approximation even more evident. To give a concrete example, we focused on the millionth weight system in the dataset, which reads $[45,74,2460,12792,17876,33173]$. Our implementation of (5.4) takes roughly $40$ hours, while the approximated version (5.5) is computed in $49$ seconds. The ratio between the two is $0.00034$. Moreover, the approximated results are very accurate: the exact ones are $(h^{1,1}=10718,\,h^{1,2}=0,\,h^{1,3}=986,\,h^{2,2}=46860)$, while the approximated ones read (10683, 0, 986, 46500)10683 098646500(10683,\,0,\,986,\,46500)( 10683 , 0 , 986 , 46500 ) [^19] To make this comparison, we had to match our conventions with the ones used for the existing datasets; this is discussed in the next section, after equation (6.2)..
Regarding the precision of the approximation, we illustrate it in Figure 3(a), with a similar plot to the one used for three-folds. We omit $h^{1,2}$ from the picture because it can take the value of zero, making the percentage error not well-defined. We sampled randomly and uniformly $20\%$ of the dataset, and then divided it into groups of $2000$ samples. As before, all samples were ordered according to the sum of weights, then dividing into the groups, as shown by the gray points. We observe once again that the percentage error gets smaller as the weights become larger, reaching roughly $1\%$ for the samples with largest weights, for all three Hodge numbers.
@@
The plot in Figure 3(b), on the other hand, shows the comparison between the computational resources employed by the exact expression from the Landau-Ginzburg model and by our approximation. It is evident from the example just discussed that systems with large weights necessitate a very long computation time for the full formula (5.4). Thus, we collected data within the first half of the dataset, and then extrapolated our findings to the second half, containing very large weights. The red curve provides a good interpolation of the data, and it turns out to give very accurate predictions as well. This can be confirmed by plugging $w_{tot}=66420$, which corresponds to the example weight system discussed above, into the expression for the best fit function. The result is $t_{app}/t=0.00035$, which is remarkably close to the actual ratio ($0.00034$) obtained from the explicit computation.
@@ 6 Higher Weight systems
The approximation has been tested on both three-folds and four-folds. These are the only two existing datasets of Calabi-Yau manifolds built as hypersurfaces in weighted projective spaces. Here, we present the first efforts towards the understanding of these spaces in higher dimensions, by generating a partial dataset of candidate transverse weight systems of 7 weights, and the respective Calabi-Yau five-folds’ Hodge numbers. We discuss how the approximation can be used to quickly extract information about such a dataset, and we additionally generate a first partial dataset of candidate transverse weight systems of 8 weights, then using (5.5) to construct an approximated list of six-folds’ Hodge numbers.
@@ 6.1 Calabi-Yau Five-folds
Calabi-Yau five-folds appear in a number of dimensional reductions in the literature. For instance, it was found that M-theory compactified on a Calabi-Yau five-fold results in an exotic $\mathcal{N}=2$ supersymmetric quantum mechanics Haupt:2008nu . Moreover, Calabi-Yau five-folds play a role in F-theory, where upon compactification, they provide a way to systematically construct $N=(0,2)$ CFTs Schafer-Nameki:2016cfr ; Tian:2020yex , which may lead to their classification. Therefore, an extended dataset of such Calabi-Yau manifolds would make it possible to explore the landscape of such CFTs. Additionally, in Curio:1998bv , a 3-dimensional string vacua with $\mathcal{N}=1$ supersymmetry has been found, which can be interpreted as a compactification of S-theory on a Calabi-Yau five-fold. Despite their role in the construction of low-dimensional theories, examples of five-folds have not been systematically constructed, until the recent effort in Alawadhi:2023gxa , which focuses on the CICY construction. Here, we present a second in that direction, i.e. we generate, for the first time, a subset of Calabi-Yau five-folds obtained as hypersurfaces in $\mathbb{P}^{6}_{\mathrm{w}}$. Specifically, we generate all 7-weight weight systems whose sum of weights $w_{tot}\leq 200$. To efficiently identify those which have the required property to describe a Calabi-Yau (for more details, see §2), we employ a two-step approach. We first systematically search all partitions of each sum of weights up to 200 as weight systems, extracting those which are coprime, IP [^20] Work in Skarke:1996hq showed that transverse weight systems are by necessity IP for any size weight system., and intradivisible; performed with our code functionality. Then we use the approximation as a tool for establishing the Calabi-Yau property and select all the weight systems that are well-defined with respect to (5.5), or equivalently all those such that the polynomial division $\left[\prod_{i}\frac{(uv)^{q_{i}}-uv}{1-(uv)^{q_{i}}}\right]$ gives no reminder. The candidate weight systems identified this way were then all checked with respect to the exact formula (5.4), and all turned out to be well-defined, yielding the full exact Hodge diamond. To provide confidence in the generated data, two non-trivial checks on the cohomological data were performed. First, computing the Euler number from the weights alone with (2.11), and then verifying it agrees with the identity
@@
As it can be guessed by looking at $\left\langle h^{1,1}\right\rangle$ and $\left\langle h^{1,4}\right\rangle$, the subset of spaces considered here does not show mirror symmetry. This is due to the fact that we only restricted ourselves to a small sum of weights, whose mirror-symmetric pairs lie in the large-weights regime. We expect the cohomological data in that regime to be practically inaccessible, due to the large computational times associated with (5.4). For this reason, we believe that the approximation presented in §5, which proved to be extremely accurate for large weights in the four-folds investigation, could be a key tool for attempting such a task. Moreover, we also expect the list of all possible Calabi-Yau 7-weight weight systems to be astronomical in size. Once again, the truncated formula (5.5) provides a quickly computable tight lower bound for the Hodge numbers of all those yet undiscovered manifolds.
@@ 6.2 Calabi-Yau Six-folds
While their role in physics is marginal (they could only be employed for compactifications of S-theory), Calabi-Yau six-folds have their own relevance directly within mathematics. These spaces could provide additional information about the - still very mysterious to this day - landscape of Calabi-Yau geometries, as they are the second non-trivial family of Calabi-Yau manifolds in even complex dimensions. Their construction as hypersurfaces of weighted projective spaces involves using 8-weight weight systems, which are both more numerous and effectively infeasible to run bulk computation of exact topological parameters. For these reasons, we find the truncated approximation formula to be especially pertinent, allowing computation of approximated Hodge values for all the generated candidate transverse weight systems with $w_{tot}\leq 200$. Once again, we first identify the IP intradivisible weight systems, and then select the ones which are well-defined with respect to (5.5), numbering 1482022 candidate transverse weight systems of 8 weights (accessible at this work’s respective GitHub in the same format as for the five-folds). From there the Euler number was also computed exactly with the less computationally-intensive direct formula from weights (2.11). A few examples are reported in Table 6.2.
@@
Similarly to before, the first manifold in Table 6.2 is nothing but the Calabi-Yau six-fold defined by a degree 8 polynomial in $\mathbb{P}^{7}$. We find that our results, despite coming from the truncated formula, exactly match the exact Hodge numbers, which can be found in Dumachev2015CompleteIC .
@@ 7 Summary & Outlook
While all of the above was performed using concrete analytic algorithms, statistical machine learning techniques were also applied both to the dataset of Calabi-Yau four-folds and to the partitioned set of more general weight systems. Regarding the former, a fully connected regressor network was shown to predict the cohomological Hodge data, and the Euler number, from the system weights. We found particularly good results, with $R^{2}\sim 0.91$, for $h^{1,1}$, on the whole dataset. For the other invariants, we observed very different results for systems with small weights as opposed to systems with large weights. For instance, $h^{1,3}$ and $h^{2,2}$ showed results with $R^{2}>0.90$ for the half of the dataset containing lower weights, while the accuracy dropped significantly for the other half. These three numbers provide sufficient information to determine the full Hodge diamond, however results were also reported associated with $h^{1,2}$, which had a poor performances since it is zero $48\%$ of the time, and $\chi$, which showed similar results to $h^{1,3}$ and $h^{2,2}$.
@@
Motivated by the strong performances of the neural networks, and inspired by the interpretability of the gradient saliency analysis and symbolic regression, we explored a simpler truncated version of the formula coming from the Landau-Ginzburg model used for calculating the Calabi-Yau Hodge numbers from the ambient $\mathbb{P}^{n}_{\textbf{w}}$’s weight system. This approximation drastically reduces the number of terms involved in the computation, making it easier to study analytically, and substantially faster to compute numerically. Its main features are: it provides a tight lower bound for the Hodge numbers; it is especially accurate for systems with large weights (average MAPE of $<1\%$ for the $10000$ systems with largest weights); it is dramatically faster than the exact formula (up to $10^{4}$ times quicker); it reproduces the observed linear clustering behaviour for large weights.
Finally, motivated by the speed improvements available from this approximation, transverse weight systems (satisfying the necessary intradivisible and IP properties, and well-defined with respect to both the approximation and exact Landau-Ginzburg formula) were generated for a sum of weights $w_{tot}\leq 200$, for systems of 7-weights producing Calabi-Yau five-folds. Additionally, where the exact Landau-Ginzburg formula computation time was infeasible for systems of 8-weights, a complementary dataset of candidate transverse weight systems (satisfying the necessary intradivisible and IP properties, and well-defined with respect to just the approximation) was generated, again for a sum of weights $w_{tot}\leq 200$, leading to candidate Calabi-Yau six-folds. Some preliminary analysis of this data, and the respectively computed topological properties is provided, with a thorough analysis, and its full generation for $w_{tot}>200$ left for future work.
These datasets, the respective code for analysis and ML, and an example notebook illustrating functionality to check intradivisibility, compute Euler number, and compute exact and approximated Hodge numbers of an input weight system of any size, are all available at this work’s respective GitHub repository.